# FLINT: Learning-based Flow Estimation and Temporal Interpolation for Scientific Ensemble Visualization

### v1

@@ VII Discussion and Conclusion
ACKNOWLEDGMENTS
We thank the Center for Information Technology of the University of Groningen for their support and for providing access to the Hábrók high performance computing cluster. We also thank the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for supporting this work by funding SFB 1313 (Project Number 327154368).
Appendix A Appendix
Appendix B Supplementary Material
B-A Scientific Ensembles
@@
B-B Density Interpolation and Flow Estimation
Illustrative results showcasing FLINT’s performance at various interpolation rates for both the LBS and Droplets ensembles are provided in Figs. 17 and 18. These results affirm the positive outcomes in terms of density interpolation and flow estimation achieved at the interpolation rates of 3 and 8, indicating the method’s reliability in these scenarios. However, it’s worth noting that at a very high rate of 32, although density interpolation can still maintain satisfactory quality, there becomes a noticeable degradation in the quality of flow estimation, which is not surprising due to the large variation in structure between the two fields.
{old}
B-C Timestep Selection
B-D Comparison Against Baselines
GT
FLINT
RIFE
In the context of the Droplets ensemble, the results presented in Fig. 13 provide evidence of FLINT’s superior performance compared to linear interpolation. This observation holds true across a spectrum of interpolation rates, according to our expectations. FLINT also consistently surpasses the RIFE {old} baselinemethod in terms of both PSNR and, previously shown in the paper, LPIPS metrics.
Figure 13: Droplets ensemble: comparison of PSNR scores of FLINT, RIFE, and linear interpolation at various interpolation rates.
B-E 5Jets, Transfer Functions
(a) TF for density field.
(b) TF for flow field in $x$ direction.
(c) TF for flow field in $y$ direction.
(d) TF for flow field in $z$ direction.
Figure 15: TF for density and $x$, $y$, $z$ components of the flow field, 5Jets.
B-F Nyx, Transfer Functions
(a) TF for density field.
(b) TF for flow field in $x$ direction.
(c) TF for flow field in $y$ direction.
(d) TF for flow field in $z$ direction.
Figure 16: TF for density and $x$, $y$, $z$ components of the flow field, Nyx.
@@ B-F Nyx, Transfer Functions
(a) Density interpolation 8$\times$: comparing FLINT predictions against GT, from $t=4606$. Rows from top to bottom: GT, interpolation, and difference.
@@
(b) Flow estimation 8$\times$: comparing FLINT predictions against GT, from $t=4606$. Rows from top to bottom: GT, flow estimation, and difference.
@@
(c) Density interpolation 32$\times$: comparing FLINT predictions against GT, from $t=4606$. Rows from top to bottom: GT, interpolation, and difference.
@@
(d) Flow estimation 32$\times$: comparing FLINT predictions against GT, from $t=4606$. Rows from top to bottom: GT, flow estimation, and difference.
@@
Figure 18: Comparison of FLINT predictions at 8$\times$ and 32$\times$ interpolation. For both density interpolation and flow estimation, results are compared against the ground truth (GT). Each row shows: GT, FLINT predictions, and the difference.

### v2

@@ VII Discussion and Conclusion
A Appendix
Supplementary Material
A-A Scientific Ensembles
@@
A-B Density Interpolation and Flow Estimation
Illustrative results showcasing FLINT’s performance at various interpolation rates for both the LBS and Nyx ensembles are provided in Figs. 1112, and 13. These results affirm the positive outcomes in terms of density interpolation and flow estimation achieved at the interpolation rates of 3 and 8, indicating the method’s reliability in these scenarios. However, it’s worth noting that at a very high rate of 32, as can be seen in Figs. 12(d) and 13, although density interpolation can still maintain satisfactory quality, there becomes a noticeable degradation in the quality of flow estimation, which is not surprising due to the large variation in structure between the two fields. In Fig. 12(d), we observe significantly larger errors for flow estimation compared to lower interpolation rates, highlighting the challenges posed by such high rates. Similarly, in Fig. 13, while the flow retains its overall structure, the accuracy in finer details, particularly around the circular swirling patterns, is noticeably reduced.
@@ A-B Density Interpolation and Flow Estimation
(a) Density interpolation 8$\times$: comparing FLINT estimations against GT, from $t=4606$. Rows from top to bottom: GT, interpolation, and difference.
@@
(b) Flow estimation 8$\times$: comparing FLINT estimations against GT, from $t=4606$. Rows from top to bottom: GT, flow estimation, and difference.
@@
(c) Density interpolation 32$\times$: comparing FLINT estimations against GT, from $t=4606$. Rows from top to bottom: GT, interpolation, and difference.
@@
(d) Flow estimation 32$\times$: comparing FLINT estimations against GT, from $t=4606$. Rows from top to bottom: GT, flow estimation, and difference.
@@
Figure 12: LBS ensemble: Comparison of FLINT estimations at 8$\times$ and 32$\times$ interpolation. For both density interpolation and flow estimation, results are compared against the ground truth (GT). Each row shows: GT, FLINT estimations, and the difference.
Dens GT
Dens FLINT
Flow GT
Flow FLINT
Figure 13: Nyx: FLINT flow field estimation and temporal density interpolation during inference, 32$\times$. From top to bottom, the rows show GT density, FLINT interpolated density, GT flow, and FLINT flow estimation. 3D rendering was used for the density and flow field visualization ( colors representing $x$, $y$, and $z$ flow directions respectively).
{old}
A-C Timestep Selection
A-D Comparison Against Baselines
GT
FLINT
RIFE
In the context of the Droplets ensemble, the results presented in Fig. 16 provide evidence of FLINT’s superior performance compared to linear interpolation. This observation holds true across a spectrum of interpolation rates, according to our expectations. FLINT also consistently surpasses the RIFE method in terms of both PSNR and, previously shown in the paper, LPIPS metrics.
In the context of the LBS ensemble, the results presented in Fig. 17 demonstrate FLINT’s superior performance compared to both linear interpolation and the RIFE method, also when evaluated using the LPIPS metric. Across varying interpolation rates, FLINT consistently achieves lower LPIPS scores, signifying better perceptual similarity to the ground truth. These findings underscore FLINT’s ability to preserve fine details and produce visually coherent interpolations, even under challenging conditions.
Figure 16: Droplets ensemble: comparison of PSNR scores of FLINT, RIFE, and linear interpolation at various rates.
A-E Noise Handling
FLINT’s robustness to noisy data was evaluated in both 2D and 3D scientific ensembles. The 2D Droplets ensemble naturally contains noisy input data, as evidenced by the density plots in Fig. 5 top row, where background noise is clearly visible. Despite this inherent noise, FLINT successfully reconstructs density and flow fields, maintaining high accuracy and capturing the essential dynamics of the dataset.
Additionally, to further test FLINT’s robustness, we introduced random Gaussian noise into the 3D Nyx simulation ensemble. Specifically, we added noise with a mean of zero and a standard deviation of 0.025 to the density field (ranging within $[0,1]$ after normalization). These parameters were selected to simulate realistic imperfections without overwhelming the dataset, allowing us to assess FLINT’s ability to maintain accuracy under noisy conditions. The results, illustrated in Fig. 18 (second row corresponds to the GT density with added noise), demonstrate that FLINT’s performance does not significantly degrade under these conditions. The model continues to produce accurate density and flow estimations (third and fifth rows, respectively), even in the presence of added noise. These findings highlight FLINT’s capability to handle real-world imperfections in data, making it a reliable solution for noisy scientific ensembles.
Dens GT
Dens GT noisy
Dens FLINT
Flow GT
Flow FLINT
Figure 18: Nyx: FLINT flow field estimation and temporal density interpolation during inference, 5$\times$. From top to bottom, the rows show GT density, GT density with added noise, FLINT interpolated density, GT flow, and FLINT flow estimation. 3D rendering was used for the density and flow field visualization ( colors representing $x$, $y$, and $z$ flow directions respectively).
A-F Evaluation on Temperature and Flow Fields
In Fig. 19, we present qualitative results for the Nyx simulation ensemble, 5×5$\times$5 ×. The temperature interpolation demonstrates good performance, with the overall structure and spatial coherence of the field being well preserved across timesteps. For the flow fields estimated from the temperature field, some deviations are observed compared to the GT. These deviations are expected, as predicting flow from temperature is inherently more challenging due to the weaker relationship between temperature distributions and physical motion. Despite this, the results still show FLINT’s capability to generate reasonable flow estimations even from scalar fields like temperature, which are less directly correlated with movement.
Temp GT
Temp FLINT
Flow GT
Flow FLINT
Figure 19: Nyx: FLINT flow field estimation and temporal temperature field interpolation during inference, 5$\times$. From top to bottom, the rows show GT temperature, FLINT interpolated temperature, GT flow, and FLINT flow estimation. 3D rendering was used for the temperature and flow field visualization ( colors representing $x$, $y$, and $z$ flow directions respectively).
(a) TF for density field.
(b) TF for flow field in $x$ direction.
(c) TF for flow field in $y$ direction.
(d) TF for flow field in $z$ direction.
Figure 20: Transfer functions for density and $x$, $y$, $z$ components of the flow field, 5Jets.
A-G 5Jets and Nyx Transfer Functions
The transfer functions used for visualizing the density field and the $x$, $y$, $z$ components of the flow field for the 5Jets and Nyx dataset are illustrated in Figs. 20 and  21. These transfer functions were selected experimentally to ensure that the rendered volumes are the most representative of the underlying data, effectively highlighting the key structures and dynamics.
(a) TF for density field.
(b) TF for flow field in $x$ direction.
(c) TF for flow field in $y$ direction.
(d) TF for flow field in $z$ direction.
Figure 21: Transfer functions for density and $x$, $y$, $z$ components of the flow field, Nyx.